Distance between two parallel planes 2x+y+2z=8 and 4x + 2y +4z+5=0 is

Grade:6

1 Answers

Akshay
6 years ago
Distance between two parallel planes ax + by + cz + d` = 0 & ax + by + cz + d" = 0 is |d` - d"|/√(a^2 + b^2 + c^2) . :. So, 2x + y + 2z - 8 =0 & 2x + y + 2z + 5/2 = 0 , then |-8 - 5/2|/√(4+1+4) = 7/2 : please approve my answer . Thank you !!